# Review and Prospect of Algebraic Research in Equivalent Framework between Statistical Mechanics and Machine Learning Theory

### v2

@@ Abstract
Mathematical equivalence between statistical mechanics and machine learning theory has been known since the 20th century, and researches based on such equivalence have provided novel methodology in both theoretical physics and statistical learning theory. For example, algebraic approach in statistical mechanics such as operator algebra enables us to analyze phase transition phenomena mathematically.
In this paper, for theoretical physicists who are interested in artificial intelligence, we review and prospect algebraic researches in machine learning theory. If a learning machine has hierarchical structure or latent variables, then the random Hamiltonian cannot be expressed by any quadratic perturbation because it has singularities. To study an equilibrium state defined by such a singular random Hamiltonian, algebraic approach is necessary to derive asymptotic form of the free energy and the generalization error. We also introduce the most recent advance, in fact, theoretical foundation for alignment of artificial intelligence is now being constructed based on algebraic learning theory.
This paper is devoted to the memory of Professor Huzihiro Araki who is a pioneer founder of algebraic research in both statistical mechanics and quantum field theory.
@@ 1 Introduction
Statistical mechanics was founded in the 19th century and have evolved throughout the 20th and 21th centuries. Information theory was established in the 20th century [59], and formal similarity between statistical mechanics and information theory has been sometimes pointed out [38, 17, 42, 13]. If nature and principle of two fields are different, such formal similarity can be treated by common mathematical foundation.
@@
Among mathematical physics in statistical mechanics and quantum field theory, algebraic approach was indicated in the 1960s by pioneer wrok by Huzihiro Araki [8, 9, 10]. In fact, types of von Neumann algebra made by a spin system was classified, quantum field theory can be treated by axiomatic way, and mathematical equivalence of equilibrium state and variational principle was clarified.
@@
This paper consists of four sections. In the second section, we explain the well-known formal equivalence of statistical physics and machine learning theory. In subsection 2.1, structure of statistical physics is described with which physicists is familiar, however, this subsection is prepared for comparison with the following subsections. In subsection 2.2, structure of machine learning theory is introduced. Physicists understand that machine learning theory is equivalent to statistical mechanics with random Hamiltonian. In subsection 2.3, we explain how concepts in statistical mechanics correspond to ones in machine learning theory. In the third section, we review and prospect algebraic approach in machine learning theory. If a learning machine contains hierarchical structure or hidden variables, then the random Hamiltonian has singularities, resulting that algebraic approach is necessary. In subsection 3.1, the main results in algebraic approach of machine learning theory are summarized. It is shown that asymptotic form of the free energy and generalization error are represented by two birational invariants, the real log canonical threshold and singular fluctuation. In subsection 3.2, mathematical properties of real log canonical threshold are clarified and its application to estimation of the free energy is explained. In subsection 3.3, mathematical properties of singular fluctuation are studied and its application to estimation of the generalization error is examined. In subsection 3.4, the most recent researches about artificial intelligence alignment is prospected. We expect that mathematical researches about statistical physics and machine learning theory would be a foundation of the future artificial intelligence development.
@@ 2 Equivalent Framework of Statistical Mechanics and Statistical Learning Theory
In this section we summarize the well-known common framework of statistical mechanics and statistical learning theory [20, 60, 54, 36, 47, 44, 62]. Even if statistical mechanics and statistical learning theory are based on the different nature and principle, their formal structure can be studied based on the common mathematical framework. From the mathematical point of view, there is no need to distinguish them.
2.1 Statistical Mechanics of Random Hamiltonian
First, we explain the formal framework of statistical mechanics, which is well-known in statistical physicists, however, this subsection is necessary for comparing with the following subsections.
Let $J$ be a random variable defined on a finite dimensional Euclidean space and $w\in\mathbb{R}^{d}$ be an observable. In order to examine a physical system with random interaction, a function $H(w,J)$ is set as a candidate model, which is referred to as a random Hamiltonian. For a given random variable $J$, the equilibrium state with an inverse temperature $\beta>0$ is defined by a probability density function of $w$,
$\rho(w|J)=\frac{1}{Z(\beta,J)}\exp(-\beta H(w,J))\varphi(w),$ (1)
[…]
$F(\beta,J)=-\frac{1}{\beta}\log Z(\beta,J).$
The average free energy over $J$ is denoted by $\mathbb{E}_{J}[F(\beta,J)]$. It is well known in statistical mechanics that, if the average free energy is explicitly calculated as a function of $\beta$, then several important values can be derived from the average free energy. Validation of modeling a random Hamiltonian is performed by comparing the theoretical results with the experimental observation.
Example in Statistical Mechanics. Let $J=\{J_{ij}\}$ and $w=\{w_{i}\}$. For example, as a probability distribution of $J$, a normal distribution is sometimes employed. A physical model of a spin glass system [20, 60] is defined by a random Hamiltonian,
$H(w,J)$ =$=-\sum_{\{i,j\}}J_{ij}w_{i}w_{j}-\sum_{i}h_{i}w_{i}$ (2)
$\chi(\beta)=\mathbb{E}_{J}\mathbb{E}_{w}^{\beta}\left[\frac{1}{n}\sum_{i}w_{i}\right]=\frac{1}{n}\sum_{i}\frac{\partial}{\partial h_{i}}\mathbb{E}_{J}[F_{n}(\beta,J)].$
2.2 Statistical Learning Theory
Second, we explain machine learning theory. Statistical physicists can understand that machine learning theory is formally equivalent to statistical mechanics.
Let $d$, $n$, $N$ are positive finite integers. In stead of random interaction $J$, we use a notation for a random sample $X^{n}$. Let
$X^{n}\equiv\{X_{i}\;;\;i=1,2,...,n\}$
be a set of $\mathbb{R}^{N}$ valued random variables which are independently subject to a probability density function $q(x)$. A learning machine is defined by a conditional probability density function $p(x|w)$ and a probability density function $\varphi(w)$ where $x\in\mathbb{R}^{N},w\in\mathbb{R}^{d}$. In statistical learning theory, $q(x)$ is an unknown data-generating distribution, and a pair of $p(x|w)$ and $\varphi(w)$ is a candidate model which is designed by a user. There is no guarantee that a model will be appropriate for an unknown data-generating process. Therefore we need mathematical theory which holds for an arbitrary triple $(q(x),p(x|w),\varphi(w))$.
A random Hamiltonian
$H(w,X^{n})=-\sum_{i=1}^{n}\log p(X_{i}|w),$ (3)
is called a posterior distribution, where
$Z(\beta,X^{n})=\int\varphi(w)\prod_{i=1}^{n}p(X_{i}|w)^{\beta}dw.$
This definition is equal to eq.(1) because $\exp(-\beta H(w,X^{n}))=\prod_{i=1}^{n}p(X_{i}|w)^{\beta}$. The expectation and variance using $\rho(w|X^{n})$ are also denoted by $\mathbb{E}_{w}^{\beta}[\;\;]$ and $\mathbb{V}_{w}^{\beta}[\;\;]$, respectively. The inference methods $\beta=1$ and $\beta=\infty$ are called Bayesian and maximum likelihood ones, respectively. The free energy is given by
$F(\beta,X^{n})=-\frac{1}{\beta}\log Z(\beta,X^{n}).$
Then random variables $Z(1,X^{n})$ and $F(1,X^{n})$ are referred to as the marginal likelihood and the minus log marginal likelihood, both of which are important values in statistics. Also $F(1,X^{n})$ is called stochastic complexity or Bayesian code length in information theory. Note that $Z(1,X^{n})$ is a probability density function
[…]
and the average free energy is equal to the entropy and the relative entropy.
[…]
where $q(x^{n})=\prod_{i=1}^{n}q(x_{i})$. This equation shows that the smaller free energy is equivalent to the smaller relative entropy. The posterior predictive distribution is defined by the average of $p(x|w)$ in the equilibrium state,
$p(x|X^{n})=\mathbb{E}_{w}^{\beta}[p(x|w)].$
Then the generalization error is defined by the relative entropy of $q(x)$ and $p(x|X^{n})$,
$G_{n}(\beta)=\int q(x)\log\frac{q(x)}{p(x|X^{n})}dx,$
which measures a difference between the data-generating function $q(x)$ and the posterior predictive distribution $p(x|X^{n})$. The smaller generalization error shows that a pair $p(x|w)$ and $\varphi(w)$ is more appropriate for the unknown data-generating distribution. Hence, one of the most important purposes in statistical learning theory is to clarify the behavior of the generalization error $G_{n}(\beta)$ as a function of $n$. However, the integration over $q(x)$ in the definition $G_{n}(\beta)$ cannot be performed, because $q(x)$ is unknown. On the other hand, the training error
$T_{n}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{q(X_{i})}{p(X_{i}|X^{n})}$
is used to estimate the generalization error. One of the purposes of theoretical study is to derive the difference between the generalization and training errors
$\mathbb{E}_{X^{n}}[G_{n}(\beta)-T_{n}(\beta)]$
[…]
which is not equal to zero in general because $X_{i}$ and $X^{n}$ are not independent. If $\beta=1$, then
$p(X_{n+1}|X^{n})=\frac{Z(1,X^{n+1})}{Z_{n}(1,X^{n})},$
resulting that
$\mathbb{E}_{X^{n}}[G_{n}(1)]$ $=\mathbb{E}_{X^{n+1}}[F(1,X^{n+1})-F(1,X^{n})]-S,$ (5)
where
[…]
is the entropy of the data-generating distribution. The relation eq.(5) shows that the generalization error is derived from the free energy if it is given by an explicit function of $n$. In machine learning, for the validation of candidate models $p(x|w)$ and $\varphi(w)$, $F(1,X^{n})$ and $G_{n}(1)$ are examined.
Example. Let $x=(x_{1},x_{2})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ and assume that $\{(X_{1i},X_{2i});i=1,2,...n\}$ is a set of independent random variables whose probability density function is $q(x_{1},x_{2})$. For the statistical learning for $q(x_{2}|x_{1})$
$p(x|w)=\frac{q(x_{1})}{(2\pi\sigma^{2})^{N/2}}\exp\left(-\frac{1}{2\sigma^{2}}(x_{2}-f(x_{1},w_{1}))^{2}\right)$ (6)
where $f(x,w_{1})$ is a function from $\mathbb{R}^{M}$ to $\mathbb{R}^{N}$. In deep learning, a function $f(x,w)$ is defined by using a layered neural network. The parameter is $w=(w_{1},\sigma)$. Remark that, in this model, $q(x_{1})$ is not estimated.
2.3 Correspondence between Physics and Machine Learning
In this subsection, we summarize the correspondence between statistical mechanics and machine learning, which is displayed in Table 1.
In statistical mechanics, the random Hamiltonian is devised as a candidate model for studying a phenomenon which has random interaction such as spin glass systems. Sometimes an assumption is set that $J$ is subject to normal distribution, validation of which is performed by comparing theoretical values with experimental observation. In statistical mechanics, it is sometimes said that, even if a candidate model is not completely constructed, the essential property of natural phenomenon can be captured by renormalizing procedure, which is referred to as the universality of theoretical physics.
In Bayesian statistics [25, 26, 46], a pair of $p(x|w)$ and $\varphi(w)$ is a candidate model for estimating unknown information source. In statistics, it is premised that all models are wrong [16, 15]. If a pair is designed, then $X^{n}$ is a set of exchangeable random variables. For exchangeable random variables, by de Finetti-Hewitt-Savage theorem [33], there exists a functional probability distribution $Q$ such that
$q(x)$ $\sim Q,$ (7)
$X^{n}$ $\sim\prod_{i=1}^{n}q(x_{i}).$ (8)
[…]
In statistical mechanics, thermodynamical limit plays the important roles, by which a macroscopic phenomenon is derived from a microscopic stochastic dynamics. For example, the dimension of $w=\{w_{i}\}$ is made to be infinity in spin glass theory. Also in machine learning theory, thermodynamical limit can taken if the leaning machine has homogeneous structure. In fact, if the function $f(x_{1},w_{1})$ in eq.(6) is given by the inner product of $x_{1}$ and $w_{1}$,
$f(x_{1},w_{1})=(x_{1})\cdot(w_{1})$
then it is possible to study thermodynamical limit. However, non-homogeneous physical phenomena such as Bose-Einstein condensation, it is impossible to study thermodynamical limit. In this paper we study nonhomogeneous learning machines such as hierarchical and degenerate neural networks which contain complex singularities, resulting that there may not exist the thermodynamical limit. It is the problem for the future study how to examine the large scale limit of nonhomogeneous learning machines. Almost all learning machines are singular [72] and that’s good [84], which may be one of the most essential properties of brain-like inference systems.
@@ 2.3 Correspondence between Physics and Machine Learning
$F(\beta,J)$ free energy $F(\beta,X^{n})$ stochastic complexity

### v3

@@ Abstract
Mathematical equivalence between statistical mechanics and machine learning theory has been known since the 20th century, and research based on this equivalence has provided novel methodologies in both theoretical physics and statistical learning theory. It is well known that algebraic approaches in statistical mechanics such as operator algebra enable us to analyze phase transition phenomena mathematically. In this paper, we review and prospect algebraic research in machine learning theory for theoretical physicists who are interested in artificial intelligence.
If a learning machine has a hierarchical structure or latent variables, then the random Hamiltonian cannot be expressed by any quadratic perturbation because it has singularities. To study an equilibrium state defined by such a singular random Hamiltonian, algebraic approaches are necessary to derive the asymptotic form of the free energy and the generalization error.
We also introduce the most recent advance: the theoretical foundation for the alignment of artificial intelligence is now being constructed based on algebraic learning theory.
This paper is devoted to the memory of Professor Huzihiro Araki who is a pioneering founder of algebraic research in both statistical mechanics and quantum field theory.
@@ 1 Introduction
Statistical mechanics was founded in the 19th century and have evolved throughout the 20th and 21th centuries. Information theory was established in the 20th century [1], and formal similarity between statistical mechanics and information theory has been sometimes pointed out [2, 3, 4, 5]. Even if nature and principle of two fields are different, such formal similarity can be treated by common mathematical foundation.
@@
Among mathematical physics in statistical mechanics and quantum field theory, algebraic approach was indicated in the 1960s by pioneer work by Huzihiro Araki [13, 14, 15]. In fact, types of von Neumann algebra made by a spin system was classified, quantum field theory can be treated by axiomatic way, and mathematical equivalence of equilibrium state and variational principle was clarified.
@@
This paper consists of four sections. In the second section, we explain the well-known formal equivalence of statistical physics and machine learning theory. In subsection 2.1, structure of statistical physics is described in a manner familiar for physicists, however, this subsection is prepared for comparison with the following subsections. In subsection 2.2, structure of machine learning theory is introduced. Physicists understand that machine learning theory is equivalent to statistical mechanics with random Hamiltonian. In subsection 2.3, we explain how concepts in statistical mechanics correspond to ones in machine learning theory. In the third section, we review and prospect algebraic approach in machine learning theory. If a learning machine contains hierarchical structure or hidden variables, then the random Hamiltonian has singularities, resulting that algebraic approach is necessary. In subsection 3.1, the main results in algebraic approach of machine learning theory are summarized. It is shown that asymptotic form of the free energy and generalization error are represented by two birational invariants, the real log canonical threshold and singular fluctuation. In subsection 3.2, mathematical properties of real log canonical threshold are clarified and its application to estimation of the free energy is explained. In subsection 3.3, mathematical properties of singular fluctuation are studied and its application to estimation of the generalization error is examined. In subsection 3.4, the most recent researches about artificial intelligence alignment is prospected. We expect that mathematical researches about statistical physics and machine learning theory would be a foundation of the future artificial intelligence development.
@@ 2 Equivalent Framework of Statistical Mechanics and Statistical Learning Theory
In this section we summarize the well-known common framework of statistical mechanics and statistical learning theory [6, 7, 8, 9, 10, 11, 12]. Even if statistical mechanics and statistical learning theory are based on different nature and principle, their formal structure can be studied based on a common mathematical framework, hence, from the mathematical point of view, there is no need to distinguish them. The correspondence between statistical mechanics and machine learning theory is summarized in Table 1.
@@
$F(\beta,J)$ free energy $F(\beta,X^{n})$ free energy
@@
This table shows the common mathematical framework of statistical mechanics and machine learning theory. Both $H(w,J)$ and $H(w,X^{n})$ are made as candidate models, whose validations are performed by comparing theoretical and observed free energies.
2.1 Statistical Mechanics of Random Hamiltonian
First, we explain the formal framework of statistical mechanics, which is well-known for statistical physicists; this subsection is necessary for comparing with the following subsections.
Let $J$ be a random variable defined on a finite dimensional Euclidean space and $w\in\mathbb{R}^{d}$ be an observable. In order to examine a physical system with random interaction, a function $H(w,J)$ of $(w,J)$ is set as a candidate model, which is referred to as a random Hamiltonian. For a given random variable $J$, the equilibrium state with an inverse temperature $\beta>0$ is defined by a probability density function of $w$,
$\rho(w|J)=\frac{1}{Z(\beta,J)}\exp(-\beta H(w,J))\varphi(w),$ (1)
[…]
$F(\beta,J)=-\frac{1}{\beta}\log Z(\beta,J).$
The average free energy over $J$ is denoted by $\mathbb{E}_{J}[F(\beta,J)]$. It is well known in statistical mechanics that, if the average free energy is explicitly calculated as a function of $\beta$, then several important theoretical values in physics can be derived from the average free energy. Validation of modeling a random Hamiltonian is performed by comparing the theoretical results obtained from the average free energy with the experimental observation.
Example in Statistical Mechanics. Let $J=\{J_{ij}\}$ and $w=\{w_{i}\}$, where $w_{i}=\pm 1$ or some continuous numbers. For a probability distribution of $J$, a normal distribution or another one is sometimes employed. A physical model of a spin glass system [6, 7] is defined by a random Hamiltonian,
$H(w,J)$ =$=-\sum_{\{i,j\}}J_{ij}w_{i}w_{j}-\sum_{i}h_{i}w_{i}$ (2)
$M(\beta)=\mathbb{E}_{J}\mathbb{E}_{w}^{\beta}\left[\frac{1}{n}\sum_{i}w_{i}\right]=\frac{1}{n}\sum_{i}\frac{\partial}{\partial h_{i}}\mathbb{E}_{J}[F(\beta,J)].$
2.2 Statistical Learning Theory
Second, we explain machine learning theory. Statistical physicists will be able to see that machine learning theory is formally equivalent to statistical mechanics.
Let $d$, $n$, $N$ are positive finite integers. In stead of a random interaction $J$, we use a notation for a random sample $X^{n}$. Let
$X^{n}\equiv\{X_{i}\;;\;i=1,2,...,n\}$
be a set of $\mathbb{R}^{N}$ valued random variables which are independently subject to an unknown probability density function $q(x)$. A learning machine is defined by a conditional probability density function $p(x|w)$ and a probability density function $\varphi(w)$ where $x\in\mathbb{R}^{N},w\in\mathbb{R}^{d}$. In statistical learning theory, $q(x)$ is an unknown data-generating distribution, and a pair of $p(x|w)$ and $\varphi(w)$ is a candidate model which is designed by a user. There is no guarantee that a model will be appropriate for an unknown data-generating process, hence we need mathematical theory which holds for an arbitrary triple $(q(x),p(x|w),\varphi(w))$.
A random Hamiltonian
$H(w,X^{n})=-\sum_{i=1}^{n}\log p(X_{i}|w),$ (3)
$=\frac{1}{Z(\beta,X^{n})}\varphi(w)\exp(-\beta H(w,X^{n}))$ (4)
is called a posterior distribution, where
$Z(\beta,X^{n})=\int\varphi(w)\prod_{i=1}^{n}\exp(-\beta H(w,X^{n}))dw.$
This definition is equal to eq.(1). The expectation and variance using $\rho(w|X^{n})$ are also denoted by $\mathbb{E}_{w}^{\beta}[\;\;]$ and $\mathbb{V}_{w}^{\beta}[\;\;]$, respectively. The inference methods $\beta=1$ and $\beta=\infty$ are called Bayesian and maximum likelihood ones, respectively. The free energy of a learning machine is given by
$F(\beta,X^{n})=-\frac{1}{\beta}\log Z(\beta,X^{n}).$
Then random variables $Z(1,X^{n})$ and $F(1,X^{n})$ are referred to as the marginal likelihood and the minus log marginal likelihood, both of which are important values in statistics. Also $F(1,X^{n})$ is called stochastic complexity or Bayesian code length in information theory. Note that $Z(1,X^{n})$ is a probability density function of $X^{n}$ defined by a model and a prior, because
[…]
and the average free energy is equal to the sum of the entropy and the relative entropy.
[…]
where $S$ is the entropy of the unknown data-generating distriution
[…]
and $q(x^{n})=\prod_{i=1}^{n}q(x_{i})$. This equation shows that the smaller free energy is equivalent to the smaller relative entropy between $q(x^{n})$ and $Z(1,X^{n})$. The posterior predictive distribution is defined by the average of $p(x|w)$ using the posterior distribution
$p(x|X^{n})=\mathbb{E}_{w}^{\beta}[p(x|w)],$
which shows the conditional probability function of $x$ for a given $X^{n}$ using $p(x|w)$ and $\varphi(w)$. Then the generalization error is defined by the relative entropy of $q(x)$ and $p(x|X^{n})$,
$G_{n}(\beta)=\int q(x)\log\frac{q(x)}{p(x|X^{n})}dx,$
which measures a difference between the unknown data-generating distribution $q(x)$ and the posterior predictive distribution $p(x|X^{n})$. The smaller generalization error shows the smaller relative entropy of $q(x)$ and $p(x|X^{n})$, which shows the a model and a prior are more appropriate for the unknown data-generating distribution according to prediction accuracy. Hence, one of the most important purposes in statistical learning theory is to clarify the behavior of the generalization error $G_{n}(\beta)$ as a function of $n$. However, the integration over $q(x)$ in the definition $G_{n}(\beta)$ cannot be performed in real applications because $q(x)$ is unknown. On the other hand, the training error
$T_{n}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{q(X_{i})}{p(X_{i}|X^{n})}$
is used to estimate the generalization error. One of the purposes of statistical learning theory is to derive the difference between the generalization and training errors
$\mathbb{E}_{X^{n}}[G_{n}(\beta)-T_{n}(\beta)]$
[…]
which is not equal to zero in general because $X_{i}$ and $X^{n}$ are not independent. If $\beta=1$, then by the definition of the predictive distribution,
$p(X_{n+1}|X^{n})=\frac{Z(1,X^{n+1})}{Z_{n}(1,X^{n})},$
resulting that
$\mathbb{E}_{X^{n}}[G_{n}(1)]$ $=\mathbb{E}_{X^{n+1}}[F(1,X^{n+1})-F(1,X^{n})]-S.$ (6)
The relation eq.(6) shows that the generalization error for $\beta=1$ is derived from the free energy if it is given by an explicit function of $n$. In machine learning, for the validation of a candidate model and prior $p(x|w)$ and $\varphi(w)$, $F(1,X^{n})$ and $G_{n}(1)$ are examined. Although $F(1,X^{n})$ and $G_{n}(1)$ have a mathematical relation eq.(6), the pair of model and prior that minimizes $F(1,X^{n})$ is different from one that minimizes $G_{n}(1)$, in general.
Example in Machine Learning. Let $x=(x_{1},x_{2})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ and assume that $\{(X_{1i},X_{2i});i=1,2,...n\}$ is a set of independent random variables whose probability density function is $q(x_{1})q(x_{2}|x_{1})$. A learning machine is sometimes employed
$p(x_{1},x_{2}|w1,\sigma)=\frac{q(x_{1})}{(2\pi\sigma^{2})^{N/2}}\exp\left(-\frac{1}{2\sigma^{2}}(x_{2}-f(x_{1},w_{1}))^{2}\right),$ (7)
where $f(x,w_{1})$ is a function from $\mathbb{R}^{M}$ to $\mathbb{R}^{N}$ which has a parameter $w_{1}$. In deep learning, a function $f(x,w_{1})$ is defined by using a layered neural network and the parameter is $w=(w_{1},\sigma)$. Remark that, in this model, an unknown conditional probability distribution $q(x_{2}|x_{1})$ is estimated whereas $q(x_{1})$ is not estimated.
2.3 Correspondence between Statistical Mechanics and Machine Learning
In this subsection, we summarize the correspondence between statistical mechanics and machine learning, which is displayed in Table 1. There are much more relations than this table between statistical mechanics and machine learning theory. The mean field approximation in the statistical mechanics is equivalent to variational Bayesian inference in machine learning [16, 17], The Langevin equation in statistical mechanics is just equal to the steepest descent with random noise in neural network learning, and the probability distribution of the random free energy gives the foundation of design of the most powerful test in statistical hypothesis testing [18].
In statistical mechanics, the random Hamiltonian is devised as a candidate model for studying a phenomenon which has random interaction such as spin glass systems. Sometimes an assumption is set that $J$ is subject to normal distribution, validation of which is performed by comparing theoretical values with experimental observation. In statistical mechanics, even if a candidate model cannot describe microscopic interaction in detail, renormalizing enables us to understand the characteristics of macroscopic phenomena. This property is called universality in theoretical physics.
In Bayesian statistics [19, 20, 21], a pair of $p(x|w)$ and $\varphi(w)$ is a candidate model for estimating an unknown uncertainty or an unknown information source $q(x)$. In statistics, it is premised that all models are wrong [22, 23] and the more appropriate pair of $p(x|w)$ and $\varphi(w)$ depends on the sample size $n$. If a pair, $p(x|w)$ and $\varphi(w)$, is designed, then $X^{n}$ is a set of exchangeable random variables. For exchangeable random variables, by de Finetti-Hewitt-Savage theorem [24], there exists a functional probability distribution $Q$ such that
$q(x)$ $\sim Q,$ (8)
$X^{n}$ $\sim\prod_{i=1}^{n}q(x_{i}).$ (9)
[…]
In statistical mechanics, thermodynamical limit sometimes plays important roles, by which a macroscopic phenomenon is derived from a microscopic stochastic dynamics. For example, the dimension of $w=\{w_{i}\}$ is made to be infinity in spin glass theory. Also in machine learning theory, thermodynamical limit can be taken if the leaning machine has homogeneous structure. In fact, if the function $f(x_{1},w_{1})$ in eq.(7) is given by the inner product of $x_{1}$ and $w_{1}$,
$f(x_{1},w_{1})=(x_{1})\cdot(w_{1})$
then it is possible to study thermodynamical limit. However, for non-homogeneous physical phenomena such as Bose-Einstein condensation, it seems to be impossible to study thermodynamical limit. In this paper we study nonhomogeneous learning machines such as hierarchical and degenerate neural networks which contain complex singularities, resulting that there may not exist the thermodynamical limit. It would be an important problem for the future study how to examine the large scale limit of nonhomogeneous learning machines without using thermodyamical limit. In statistical learning theory, almost all learning machines are singular [26] and that’s good [27], which seems to be one of the most essential properties of brain-like inference systems.